OpenStudy (anonymous):

The table shows the result of rolling a number cube several times ________________________________ outcome 1 2 3 4 5 6 ________________________________ number of times 5 4 7 5 4 5 Outcome ocurred ________________________________ To the nearest percent, what is the experimental probability of rolling a 3 or a 6? A. 12% B. 30% C. 33% D. 40%

OpenStudy (anonymous):

Ohhh my answer is wrong neways -_- n= 30

OpenStudy (anonymous):

ok now I understand, thank you the answer is B 30%

OpenStudy (anonymous):

P(3 or 6) = P(3) +P(6) So n is the amount of times the die is rolled which is: n = 5+4+7+5+4+5 = 30 So \[ P(3) = \large \frac{ \text { amount of times die land on 3} }{ \text{ Total times die rolled}}\] So \[P(3)=\frac{7}{30} \] And\[ P(6) = \large \frac{ \text { amount of times die land on 6} }{ \text{ Total times die rolled}}\] So \[P(3)=\frac{5}{30} \] Hence P(3 or 6) = P(3) +P(6) =\( \frac{7}{30} +\frac{5}{30}= .4\) So the probability that it will land on either 3 or 6 is .4

OpenStudy (anonymous):

Why would u think that the answer is B?

OpenStudy (anonymous):

maybe, I'm confused now, but B can be the right answer

OpenStudy (anonymous):

Ok at what point did you get confused?

OpenStudy (anonymous):

btw P(3) means the probability that the die will land on 3 and P(6) means the probability that the die will land on 6

OpenStudy (anonymous):

Because I ask the same question to weegy.com and they get 30%, so I don't know.

OpenStudy (anonymous):

But I trust in openstudy

OpenStudy (anonymous):

Ok But read what I wrote. Do you follow what I did?

OpenStudy (anonymous):

I'm doing now

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

and a percentage is a number times 100 So keep that in mind for the answer :P

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so for example if the probability is .75 then we would multiply .75*100 and we would get 75%

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

so do you follow what i did?

OpenStudy (anonymous):

So what would the answer be based on my explanation?